1. Suppose Skippy is indifferent between the following bundles of Apples and Oranges:

Units of apple Units of Oranges

1 16

2 10

3 6

4 4

Graph Skippy's indifference curve. What is Skippy's MRS between each point on the curve? What happens to the MRS as he consumes more apples? Explain.
Answer #1

→ UNITS OF ORANGE ÁY:6 2,10) ox= I 8† 18841 sxT43,6) - - - - - - + +- 1 2 3 4 5 6 7 8 9 10 UNITS OF APPLE →

So the above graph represents Skippy's Indifference curve where x-axis shows apples and y-axis shows oranges

We know, MRS = Slope of Indifference curve

and Slope = ∆y/∆x

So, MRS = ∆y/∆x , where y is oranges and x is apples

  • MRS between (1,16) and (2,10) = 6/1 = 6
  • MRS between (2,10) and (3,6) = 4/1 = 4
  • MRS between (3,6) and (4,4) = 2/1 = 2

As, Units of apples is increased from 1 to 2 then 2 to 3 and then 3 to 4 results in a fall of MRS from 6 to 4 and then to 2

So,we can see as Skippy increases consumption of apples the MRS falls.

It happens due to the property of Diminishing Marginal Rate of Substitution which states that :

Rate at which one of the good can be given up in an exchange with the other good such that the individual stays at the same utility level will decreases as the unit of the good increases.